OpenStudy (anonymous):

need to find 5 solutions to the following problem.

OpenStudy (anonymous):

\[y ^{5}-81y=0\]

OpenStudy (anonymous):

y(y^4 - 81) = 0 y(y^2 + 9)(y^2 - 9) = 0 y(y^2 + 9)(y + 3)(y - 3) = 0 \[y = 0, \pm 3, \pm 3i\]